OpenStudy (anonymous):

Simplify Using identities. CosB*SinB/Sin^2B-Cos^2B

OpenStudy (anonymous):

Try to use Identities: \[2\sin(x) \cos(x) = \sin(2x)\] \[\cos^2(x) - \sin^2(x) = \cos(2x)\]

OpenStudy (anonymous):

To use this efficiently, multiply and divide by \(-2\)..

OpenStudy (anonymous):

The 2B is actually separate. B is an angle. and Sin^2 and Cos^2. After simplifying it must be equal to cotB/1-cot^2 B. And I'm unable to simplify it that far

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Yes I know that..

OpenStudy (anonymous):

You mea\[\frac{\cos(x)\sin(x)}{\sin^2(x) - \cos^2(x)}\]n this:

OpenStudy (anonymous):

I mean that is your question no?

OpenStudy (anonymous):

So multiply and divide by \(-2\) : \[\frac{-2 \cos(x) \sin(x)}{2(\cos^2(x) - \sin^2(x))} \implies \frac{-\sin(2x)}{2\cos(2x)} \implies \color{green}{\frac{-1}{2}\tan(2x)}\]

OpenStudy (anonymous):

In denominator, I have used - sign to brackets, so cosine term will come first and then sine..

OpenStudy (anonymous):

Yes is. -1/2 tan(2X) your simplified answer?

OpenStudy (anonymous):

I have just calculated it, you should check now whether it is simplified or not.. :)

OpenStudy (anonymous):

The final simplified solution is supposed to be cotB/1-cot^2 B